OpenStudy (anonymous):

how about this one 6x+5y=90 what is the y intercept and the x intercept

OpenStudy (curry):

15

OpenStudy (aroub):

if u want to find the x -intercept substitute y=0 and y-intercept substitute x=0

OpenStudy (anonymous):

x intercept (18,0) y (0,15)

OpenStudy (anonymous):

6(0)+5y=90

OpenStudy (curry):

ishaan u mixed it up its x intercept is 15

OpenStudy (anonymous):

it is simple 6x+5y=90 x/(90/6)+y/(90/5)=1 it is of the form x/a+y/b=1 a=x intercept b=y intercept a=15 b=18

OpenStudy (anonymous):

replace y for 0 to find the interception in xx and replace x for 0 to find the interception in yy.

OpenStudy (anonymous):

i hope its right

OpenStudy (anonymous):

oh yeah mixed up sorry

OpenStudy (anonymous):

thanks curry